Can someone help me find the solution of this system using the substition method?
3x + 7y = 5
-4x +y = 14

Answer by Alan3354(69443)

3x + 7y = 5
-4x +y = 14
-----------
-4x +y = 14
--> y = 4x + 14
Sub for y in the 1st equation